# Maia-2: A Unified Model for Human-AI Alignment in Chess

### v1

@@ Abstract
There are an increasing number of domains in which artificial intelligence (AI) systems both surpass human ability and accurately model human behavior. This introduces the possibility of algorithmically-informed teaching in these domains through more relatable AI partners and deeper insights into human decision-making. Critical to achieving this goal, however, is coherently modeling human behavior at various skill levels. Chess is an ideal model system for conducting research into this kind of human-AI alignment, with its rich history as a pivotal testbed for AI research, mature superhuman AI systems like AlphaZero, and precise measurements of skill via chess rating systems. Previous work in modeling human decision-making in chess uses completely independent models to capture human style at different skill levels, meaning they lack coherence in their ability to adapt to the full spectrum of human improvement and are ultimately limited in their effectiveness as AI partners and teaching tools. In this work, we propose a unified modeling approach for human–AI alignment in chess that coherently captures human style across different skill levels and directly captures how people improve. Recognizing the complex, non-linear nature of human learning, we introduce a skill-aware attention mechanism to dynamically integrate players’ strengths with encoded chess positions, enabling our model to be sensitive to evolving player skill. Our experimental results demonstrate that this unified framework significantly enhances the alignment between AI and human players across a diverse range of expertise levels, paving the way for deeper insights into human decision-making and AI-guided teaching tools. Maia-2 implementation is available here
@@ 2 Related Work
2.1 Chess and AI
This paper draws on the long history of chess at the forefront of AI research [2, 3, 4, 5]. We engage with 3 distinct approaches to building chess AI: heuristic [6], learned [7], and textual [8].
Heuristic search. The original approach to computer chess was heuristics-based [4, 9]. This method was famously used by IBM’s Deep Blue to defeat Garry Kasparov [2] and is currently used by Stockfish [6], one of the strongest chess engines in the world.
Learned search. Alpha(Zero) Go [7, 10] is a set of neural networks that learn to play Go with methods that generalized to other games, including chess, with AlphaZero [10]. The convolutional neural network stack they introduced is what Maia-2’s is based on.
Chess as text. Large language models [11, 12, 13] have recently been found to perform well on tasks that the models were not explicitly trained on [14], including playing chess without fine-tuning  [15, 16, 17]. This has lead to chess knowledge being one of the tested features in BIG-Bench [18], a popular LLM evaluation suite. Additionally, fine-tuning a language model can lead to systems that not only play chess, but can also generate comments, describe positions, and create other simple analyses of a game [19, 20, 8].
2.2 Human-AI Alignment in Chess
Building a chess engine that can defeat any human has been a solved problem for over 20 years. This has led to a new research agenda in extracting useful knowledge from these superhuman systems. A direct way of doing this is to probe an AI chess engine in a human representation space. Without any prior human knowledge or guidance, evidence of human chess concepts learned by AlphaZero is found and measured by linear probes [21]. Going further, AlphaZero also encodes knowledge that extends beyond existing human knowledge but is ultimately learnable by humans [22].Another direction was the creation of a ‘behavioral stylometry’ model that can identify chess players from the moves they play [23].
Human-like chess AI. An alternative approach to creating systems that can act as guides to humans is demonstrated by Maia [1, 24], in which a model is trained to predict the next move a human at a given skill level will play, instead of optimizing for winning the game. In addition to predicting human actions, the models have been fine-tuned to predict a given player’s actions [24]. This method suggests that players have a distinct and measurable style that can be learned by observing a sufficient number of games. The prediction accuracy can be improved via a reinforcement learning-style search [25].
@@ 3 Methodology
We propose a unified model architecture to capture human decision-making in chess across a broad spectrum of skill levels. Since this model builds upon the previous Maia move-matching models, we call it Maia-2. As shown in Figure 1, Maia-2 first encodes active and opponent skill levels and the chess positions, respectively. Then the encoded skill levels and positions are fused using our skill-aware attention with channel-wise patching architecture. The fused representations are then used for move prediction (policy head), auxiliary information prediction (info head), and game outcome prediction (value head). We now discuss each of these components in detail.
@@ 3.3 Bridging Skill Levels and Positions
Channel-wise patching. In contrast to the area-wise patching approach in Vision Transformers (ViTs) [28], we employ channel-wise patching. Each channel is flattened and linearly transformed, regarding the number of channels in $P_{\text{encoded}}$, i.e., $C_{\text{patch}}$, as the sequence length.
$P_{\text{patched}}=\text{Patching}(P_{\text{encoded}})\in\mathbb{R}^{C_{\text{patch}}\times 64},$
$P=P_{\text{patched}}\mathbf{W}+\mathbf{b}\in\mathbb{R}^{C_{\text{patch}}\times d_{\text{att}}},$
@@
where $\mathbf{W}\in\mathbb{R}^{64\times d_{\text{att}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{\text{att}}}$ denote the parameters of the linear projection from the patching dimension to the hidden dimension of the skill-aware attention blocks $d_{\text{att}}$. This is particularly suitable for patching encoded chess positions as inputs to Transformer-like architectures, where channels are essentially feature maps that represent different learned latent concepts. These concepts in feature maps are then interactively selected and aggregated considering skill levels via skill-aware attention.
Skill-aware Attention. Given position representations $P_{\text{patched}}$ and skill level representations $\mathbf{e}_{a}$ and $\mathbf{e}_{o}$, our proposed skill-aware multi-head self-attention is computed as follows.
For each head $k$, we learn weight matrices $\mathbf{W}^{Q}_{k}\in\mathbb{R}^{d_{\text{att}}\times d_{h}}$, $\mathbf{W}^{K}_{k}\in\mathbb{R}^{d_{\text{att}}\times d_{h}}$, and $\mathbf{W}^{V}_{k}\in\mathbb{R}^{d_{\text{att}}\times d_{h}}$, where $d_{h}$ denote the dimension of each head. The queries $Q_{k}$, keys $K_{k}$, and values $V_{k}$ for each head are computed as:
$Q_{k}=P_{\text{patched}}\mathbf{W}^{Q}_{k},\quad K_{k}=P_{\text{patched}}\mathbf{W}^{K}_{k},\quad V_{k}=P_{\text{patched}}\mathbf{W}^{V}_{k}$
@@
In order to fuse player skill levels and chess positions progressively and interactively, we inject skill level embeddings into queries within the multi-head self-attention: $Q^{*}_{k}=Q_{k}+(\mathbf{e}_{a}\oplus\mathbf{e}_{o})\mathbf{W}^{*},$ where $\mathbf{W}^{*}\in\mathcal{R}^{2d_{s}\times d_{h}}$ denotes the weight matrix for feature transformation to the query space, and $\oplus$ is the concatenation operator. We choose to incorporate skill levels in queries because queries directly influence how attention is distributed across patched channels. Using skill-aware queries $Q^{*}_{k}$, the attention mechanism can adjust its focus to reflect the strategic considerations and positional understanding of players at different skill levels. This adjustment allows Maia-2 to adaptively prioritize features of the positions that are more relevant to the skill levels involved, enhancing the model’s contextual sensitivity. The skill-aware scaled dot-product attention for each head is thus defined as:
$h_{k}=\text{softmax}\left(\frac{Q^{*}_{k}K_{k}^{T}}{\sqrt{d_{k}}}\right)V_{k}$
The outputs of all heads $h_{1},{h}_{2},\ldots,{h}_{h}$ are concatenated and then linearly transformed: $P_{\text{att}}=\sigma(({h}_{1}\oplus{h}_{2}\oplus\ldots\oplus{h}_{h})\mathbf{W}^{O})$, where $\mathbf{W}^{O}\in\mathbb{R}^{hd_{h}\times d_{\text{att}}}$ denote the weight matrix for multi-head attention and $\sigma(\cdot)$ denotes the activation function. We apply the vanilla ViT’s feed-forward network and add & norm components upon $P_{\text{att}}$ to obtain the output of each skill-aware attention block $P_{\text{out}}$. In Maia-2, we employ a sequence of skill-aware attention blocks to progressively fuse skill levels and positions. Specifically, the output $P_{\text{out}}$ for the previous block is fed into the next block as the input. We denote the final output after $K_{\text{Att}}$ blocks as $P$. This procedure enables the model to refine its understanding and interpretation of the positions with each successive block.
@@ 3.4 Model Training
Infusing auxiliary information. To enhance the model’s understanding of the game state, we inject auxiliary information as labels, including legal moves represented by multi-hot vectors and human move information: one-hot vectors of which piece is moved, which piece is captured (if any), the move’s originating square, the move’s destination square, and whether or not the move will deliver a check. For nuanced moves like castling, we ensure both the king’s and rook’s moves are accurately represented with 2-hot vectors. These segments are then concatenated into a comprehensive multi-hot vector to be used as labels for classification, serving a dual purpose: 1) It offers a more granular understanding of human moves by providing detailed context beyond just the move indices produced by the policy head labels, enriching the model’s insight of player decisions; and 2) It ensures the model also learns about objective (i.e. chess-specific as opposed to behavioral) knowledge in chess, which is essential for developing a comprehensive understanding of both human moves and the fundamental mechanics of the game.
Data filtering. Online chess platforms feature a variety of game types, including blitz, rapid, and classical, each representing games played at different time controls (amount of time given to each player for the whole game). We use data from Lichess, a well-known large open-source chess platform, and its open database. In Lichess, since each game type is given a separate rating, ratings across different game types are not comparable (e.g. a rating of 1800 in “Rapid” is significantly weaker than a rating of 1800 in “Blitz” on Lichess). Previous work [1, 24] mixes player ratings across these game types together for training and evaluation. Instead of mixing data across game types, we focus on Rapid games only, which are medium-length games that lie between the fast-paced decisions of “Blitz” games and the slower, more strategic considerations of “Classical” games. “Blitz” and “Bullet” games, characterized by their quick pace, are composed of many decisions made under time pressure, introducing randomness that may not accurately reflect player intentions and skills. In contrast, Classical games are played less frequently, leading to data scarcity. Rapid chess is an ideal compromise between quality of play and quantity of data. In addition, we follow the procedures in [1] to filter valid positions within each game (more details can be found in Appendix Section 7).
@@ 4 Results
We empirically evaluate Maia-2 along two key dimensions: move prediction accuracy, how well it can predict human moves at varying skill levels, and move prediction coherence, how aligned its predictions are across skill levels.
4.1 Experimental Setup
Maia-2. We train Maia-2 on Lichess games played between Jan 2013 and Nov 2023 (inclusive), with the exception of Dec 2019, since that is the month used for testing in the original Maia paper (and we also test on this month for consistency) [1]. After game filtering and balancing, we end up with a training set of 169M games (9.1B positions) played between May 2018 and Nov 2023 (inclusive).
@@ 4.1 Experimental Setup
Model comparison. We compare Maia-2 with several baselines. The first is Stockfish [6], the strongest chess engine in the world at the time of writing. Since the most common method of using Stockfish to play in a “human-like” way is to limit its search depth, we test it at various search depths. The second is Leela, an open-source counterpart to AlphaZero [10], which we also test at various strengths, since chess commentators have anecdotally observed that its decisions are more reminiscent of human-like play at lower strengths. Finally, our main model comparison will be with Maia (which we will refer to as “Maia-1” to avoid confusion with our model), the state-of-the-art model for human-like chess play [1]. Maia-1 is actually a set of 9 separate models, each trained on a different set of players at different skill levels from 1100 to 1900. Maia-1100 models the weaker players, Maia-1500 the intermediate players, and Maia-1900 the higher-skill players.
Evaluation Datasets. We evaluate using three main test sets. To enhance fair comparison with baseline models, we use the benchmarking Maia-1 Testset  [1] for performance comparisons where both players have identical skill levels. We report the results on Maia-1 Testset by grouping players into three categories: Skilled (Rapid rating up to 1600, which slightly exceeds the initial rating of 1500), Advanced (Rapid rating between 1600 and 2000), and Master (Rapid rating over 2000, roughly comprising the top 10% of players[29]). In addition, we aim to evaluate move prediction across diverse skill combinations, which Maia-1 Testset excludes. Therefore, we construct the Cross-skill Testset, a new testing dataset from Lichess games played in Dec 2023, ensuring that positions under each skill level combination have sufficient data to be statistically reliable. Detailed statistics of this dataset are summarized in Appendix Table 8. Finally, we construct Grounded Testset by extracting 450,000 positions from Dec 2023 games in Lichess Database where Stockfish evaluations are available. Such recorded evaluations serve as grounded facts to measure move quality.
Table 1: Move prediction accuracy on the Maia-1 Testset. Skilled, Advanced, and Master are grouped according to Section 4.1 and Avg denotes macro-averaged results.
Skilled Advanced Master Avg
Stockfish 3 36.22 38.25 40.71 38.39
Stockfish 7 35.66 38.08 42.25 38.66
Stockfish 9 36.00 38.78 43.26 39.35
Stockfish 11 36.38 39.33 43.71 39.81
Stockfish 15 36.86 39.83 44.61 40.43
Leela 1500 40.46 44.45 48.69 44.53
Leela 2200 39.79 43.97 47.11 43.62
Leela 3200 39.97 44.29 47.75 44.00
Leela 3700 40.47 44.65 48.12 44.41
Maia 1100 51.48 49.13 45.85 48.82
Maia 1500 50.79 52.61 50.76 51.39
Maia 1900 48.51 52.26 53.20 51.32
@@ 4.2 Move Prediction Accuracy
In Table 1, we show the top-1 move prediction accuracy of all models across all groups of players on the Maia-1 Testset.
Maia-2. Maia-2 demonstrates strong and consistent performance across all skill levels, surpassing all baselines. Specifically, despite Maia-1 models being specifically trained to mimic chess moves by players at specific skill levels, Maia-2 emerges as a unified one-for-all model that is consistently effective across the entire spectrum of chess skills. The largest improvement is on Advanced players, where Maia-2 gains 1.5 percentage points over the nearest competitor (Maia 1500). Recall that human move prediction has a large amount of intrinsic noise: different players of equal skill facing the same position may make different points, and indeed, even the same player facing the same position may vary in their decision-making. When averaging across skill levels, Maia-2 outperforms all other models by almost 2 full percentage points in overall accuracy.
Baseline models. Both Maia-2 and Maia-1 significantly outperform Stockfish and Leela, typically by 5–15 percentage points. Note that Stockfish and Leela aim to play optimal chess (as most humans do too), and only “predict” human moves when their approximations to optimality happen to overlap with those of human players. However, we compare to these traditional chess engines because besides Maia-1, there are still the default method of creating “human-like” AI agents. The accuracy gap between Maia architectures and traditional chess engines demonstrates the necessity of developing specialized models to mimic human chess moves.
@@
[…]
Move prediction perplexity. While top-1 accuracy gains are important, they may overshadow larger improvements in prediction quality. To account for this, we also measure the perplexity of move predictions , which reflects the model’s confidence in its predictions. A lower perplexity indicates the model is more confident and accurate in human move prediction, as it corresponds to a higher likelihood (lower log-likelihood) of the correct human move. As shown in Table 2, Maia-2 consistently yields substantially lower perplexity in all groups of skill levels compared to Maia-1. In particular, Maia-2 significantly outperforms Maia-1 in Advanced and Master moves with relatively large margins, demonstrating the effectiveness of our unified modeling approach across diverse skill levels.
@@
Adaptive move predictions. We now evaluate Maia-2’s ability to predict compare across diverse skill combinations using the Cross-skill Testset.
As shown in Figure 2, Maia-2 consistently outperforms both Maia 1100 and Maia 1900 in almost all combinations of active player and opponent player skill levels. In particular, although Maia 1100 and Maia 1900 demonstrate competent performance within their respective domains of expertise, their predictive accuracy decreases substantially outside of these targeted skill levels. This is because Maia-1 models are static and cannot respond to varied skill levels and adjust their predictions accordingly. In contrast, our proposed unified modeling approach with skill-aware attention enables Maia-2 to adapt its predictions to account for the skill levels of both the active player and the opponent player, so that varying skill level configurations correspondingly can result in better aligned human move predictions. More results on comparisons with other Maia-1 versions can be found in Figure 7 in the Appendix.
Move quality. One of our key motivations for creating a unified model of human chess behavior is to guide the development of future algorithmic learning tools. As such, understanding the mistakes that people make is of fundamental interest. Can Maia-2 predict mistakes better than Maia-1? Figure 11 in the Appendix shows the move prediction accuracy on the Grounded Testset as a function of move quality, measured by win-rate loss, which is calculated following the same procedures as prior studies [1, 24]. All models generally decrease in their ability to predict worse moves, since humans are generally trying to avoid mistakes, and high-quality moves are more certain whereas lower-quality moves can be more random and thus hard to predict. Nevertheless, Maia-2 outperforms all versions of Maia across most of the move quality range, demonstrating the effectiveness of our unified modeling approach for human move prediction.
We are also interested in how certain the models are about their predictions of various move qualities. Figure 3(top) shows the probabilities that Maia-2 (x-axis) and Maia-1 (y-axis) attribute to the moves people actually played in Grounded Testset. The concentration of points in the lower right quadrant (closer to P(Maia-2) = 1 and P(Maia 1900) = 0) suggests that Maia-2 assigns a higher probability to the true move than Maia-1 does, indicating superior predictive performance. Conversely, the less dense upper left quadrant indicates fewer instances where Maia 1900 outperforms Maia-2. Remarkably, while this consistenly occurs across all move qualities, the distinction is more pronounced for Blunders and Errors compared to Optimal moves. Additionally, the bottom row of Figure 3 shows the log odds ratio between $P(x,y)$ and $P(y,x)$ in the top row. The abundance of blue points below the diagonal indicate that Maia-2 is almost always more confident in the correct move than Maia-1 is, indicating an across-the-board improvement in move prediction. Maia-2 offers superior and more confident prediction across diverse move qualities.
Figure 3: (Top) Joint probability assigned to human moves played by Maia-2 ($x$) and Maia 1900 ($y$), split by move quality. Blunders (left) reduce the expected win-rate by $\geq$ 10%, Errors (middle) by 5–10%, and Optimal (right) by $\leq 0\%$. (Bottom) Log odds ratio of $p(x,y)$ and $p(y,x)$ from top.
[…]

### v2

@@ Abstract
There are an increasing number of domains in which artificial intelligence (AI) systems both surpass human ability and accurately model human behavior. This introduces the possibility of algorithmically-informed teaching in these domains through more relatable AI partners and deeper insights into human decision-making. Critical to achieving this goal, however, is coherently modeling human behavior at various skill levels. Chess is an ideal model system for conducting research into this kind of human-AI alignment, with its rich history as a pivotal testbed for AI research, mature superhuman AI systems like AlphaZero, and precise measurements of skill via chess rating systems. Previous work in modeling human decision-making in chess uses completely independent models to capture human style at different skill levels, meaning they lack coherence in their ability to adapt to the full spectrum of human improvement and are ultimately limited in their effectiveness as AI partners and teaching tools. In this work, we propose a unified modeling approach for human-AI alignment in chess that coherently captures human style across different skill levels and directly captures how people improve. Recognizing the complex, non-linear nature of human learning, we introduce a skill-aware attention mechanism to dynamically integrate players’ strengths with encoded chess positions, enabling our model to be sensitive to evolving player skill. Our experimental results demonstrate that this unified framework significantly enhances the alignment between AI and human players across a diverse range of expertise levels, paving the way for deeper insights into human decision-making and AI-guided teaching tools. Our implementation is available here.
@@ 2 Related Work
Chess and AI. This paper draws on the long history of chess at the forefront of AI research [2, 3, 4, 5]. We engage with 3 distinct approaches to building chess AI: heuristic [6], learned [7], and textual [8]. Heuristic search: The original approach to computer chess was heuristics-based [4, 9]. This method was famously used by IBM’s Deep Blue to defeat Garry Kasparov [2] and is currently used by Stockfish [6], one of the strongest chess engines in the world. Learned search: Alpha(Zero) Go [7, 10] is a set of neural networks that learn to play Go with methods that generalized to other games, including chess, with AlphaZero [10]. Chess AI with learned search is also extended to multi-agent systems [11], where diverse AI systems can outperform a single AI in challenging tasks such as chess. Chess as text: Large language models [12, 13, 14] have recently been found to perform well on tasks that the models were not explicitly trained on [15], including playing chess without fine-tuning [16, 17, 18]. This has lead to chess knowledge being one of the tested features in BIG-Bench [19], a popular LLM evaluation suite. Additionally, fine-tuning a language model can lead to systems that not only play chess, but can also generate comments, describe positions, and create other simple analyses of a game [20, 21, 8].
Human-AI Alignment in Chess. Building a chess engine that can defeat any human has been a solved problem for over 20 years. This has led to a new research agenda in extracting useful knowledge from these superhuman systems. A direct way of doing this is to probe an AI chess engine in a human representation space. Without any prior human knowledge or guidance, evidence of human chess concepts learned by AlphaZero is found and measured by linear probes [22]. Going further, AlphaZero also encodes knowledge that extends beyond existing human knowledge but is ultimately learnable by humans [23].Another direction was the creation of a ‘behavioral stylometry’ model that can identify chess players from the moves they play [24]. An alternative approach to creating systems that can act as guides to humans is demonstrated by Maia [1, 25], in which a model is trained to predict the next move a human will play, instead of optimizing for winning the game. In addition to predicting human actions the models have been fine-tuned to predict a given player’s actions [25]. The prediction accuracy can be improved via a reinforcement learning-style search [26].
@@ 3 Methodology
We propose a unified model architecture to capture human decision-making in chess across a broad spectrum of skill levels. Since this model builds upon the previous Maia move-matching models, we call it Maia-2. As shown in Figure 1, Maia-2 first encodes active and opponent skill levels and the chess positions, respectively. Then the encoded skill levels and positions are fused using our skill-aware attention with channel-wise patching architecture. The fused representations are then used for move prediction (policy head), auxiliary information prediction (auxiliary head), and game outcome prediction (value head). We now discuss each of these components in detail.
@@ 3.3 Bridging Skill Levels and Positions
Channel-wise patching. In contrast to the area-wise patching approach in Vision Transformers (ViTs) [29], we employ channel-wise patching. Each channel is flattened and linearly transformed, regarding the number of channels in $P_{\text{encoded}}$, i.e., $C_{\text{patch}}$, as the sequence length: $P_{\text{patched}}=\text{Patching}(P_{\text{encoded}})\in\mathbb{R}^{C_{\text{patch}}\times 64}$, $P=P_{\text{patched}}\mathbf{W}+\mathbf{b}\in\mathbb{R}^{C_{\text{patch}}\times d_{\text{att}}}$, where $\mathbf{W}\in\mathbb{R}^{64\times d_{\text{att}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{\text{att}}}$ denote the parameters of the linear projection from the patching dimension to the hidden dimension of the skill-aware attention blocks $d_{\text{att}}$. This is particularly suitable for patching encoded chess positions as inputs to Transformer-like architectures, where channels are essentially feature maps that represent different learned latent concepts. These concepts in feature maps are then interactively selected and aggregated considering skill levels via skill-aware attention.
@@
Skill-aware Attention. Given position representations $P_{\text{patched}}$ and skill level representations $\mathbf{e}_{a}$ and $\mathbf{e}_{o}$, our proposed skill-aware multi-head self-attention is computed as follows. For each head $k$, we learn weight matrices $\mathbf{W}^{Q}_{k}\in\mathbb{R}^{d_{\text{att}}\times d_{h}}$, $\mathbf{W}^{K}_{k}\in\mathbb{R}^{d_{\text{att}}\times d_{h}}$, and $\mathbf{W}^{V}_{k}\in\mathbb{R}^{d_{\text{att}}\times d_{h}}$, where $d_{h}$ denote the dimension of each head. The queries $Q_{k}$, keys $K_{k}$, and values $V_{k}$ for each head are computed as: $Q_{k}=P_{\text{patched}}\mathbf{W}^{Q}_{k}$, $K_{k}=P_{\text{patched}}\mathbf{W}^{K}_{k}$, $V_{k}=P_{\text{patched}}\mathbf{W}^{V}_{k}$. In order to fuse player skill levels and chess positions progressively and interactively, we inject skill level embeddings into queries within the multi-head self-attention: $Q^{*}_{k}=Q_{k}+(\mathbf{e}_{a}\oplus\mathbf{e}_{o})\mathbf{W}^{*},$ where $\mathbf{W}^{*}\in\mathcal{R}^{2d_{s}\times d_{h}}$ denotes the weight matrix for feature transformation to the query space, and $\oplus$ is the concatenation operator. We choose to incorporate skill levels in queries because queries directly influence how attention is distributed across patched channels. Using skill-aware queries $Q^{*}_{k}$, the attention mechanism can adjust its focus to reflect the strategic considerations and positional understanding of players at different skill levels. This adjustment allows Maia-2 to adaptively prioritize features of the positions that are more relevant to the skill levels involved, enhancing the model’s contextual sensitivity. The skill-aware scaled dot-product attention for each head is thus defined as: $h_{k}=\text{softmax}\left(\frac{Q^{*}_{k}K_{k}^{T}}{\sqrt{d_{k}}}\right)V_{k}$. The outputs of all heads $h_{1},{h}_{2},\ldots,{h}_{h}$ are concatenated and then linearly transformed: $P_{\text{att}}=\sigma(({h}_{1}\oplus{h}_{2}\oplus\ldots\oplus{h}_{h})\mathbf{W}^{O})$, where $\mathbf{W}^{O}\in\mathbb{R}^{hd_{h}\times d_{\text{att}}}$ denote the weight matrix for multi-head attention and $\sigma(\cdot)$ denotes the activation function. We apply the vanilla ViT’s feed-forward network and add & norm components upon $P_{\text{att}}$ to obtain the output of each skill-aware attention block $P_{\text{out}}$. In Maia-2, we employ a sequence of skill-aware attention blocks to progressively fuse skill levels and positions. Specifically, the output $P_{\text{out}}$ for the previous block is fed into the next block as the input. We denote the final output after $K_{\text{Att}}$ blocks as $P$. This procedure enables the model to refine its understanding and interpretation of the positions with each successive block.
@@ 3.4 Model Training
Infusing auxiliary information. To enhance the model’s understanding of the game state, we inject auxiliary information as labels, including legal moves represented by multi-hot vectors and human move information: one-hot vectors of which piece is moved, which piece is captured (if any), the move’s originating square, the move’s destination square, and whether or not the move will deliver a check. These segments are used as labels for classification, serving a dual purpose: 1) It offers a more granular understanding of human moves by providing detailed context beyond just the move indices produced by the policy head labels, enriching the model’s insight of player decisions; and 2) It ensures the model also learns about objective (i.e. chess-specific as opposed to behavioral) knowledge in chess, which is essential for developing a comprehensive understanding of both human moves and the fundamental mechanics of the game.
Data balancing and filtering. Chess games between players of significantly different skill levels are relatively rare but help us understand how players of lower skill levels approach games against far stronger opponents and vice versa. While previous work has ignored these games completely, they play a central role in our approach. Since games between players of similar skill levels vastly outnumber more uneven matchups, we use a data balancing strategy to effectively train our unified model for aligning players across all skill levels, in which games between players of different skill levels are over-sampled. Online chess platforms feature a variety of game types, including blitz, rapid, and classical, each representing games played at different time controls (amount of time given to each player for the whole game). We focus on Rapid games, which are medium-length games that lie between the fast-paced decisions of “Blitz” games and the slower, more strategic considerations of “Classical” games. In addition, we follow the procedures in [1] to filter valid positions within each game. More details about data balancing and filtering are available in Appendix 7.
@@ 4 Results
Table 1: Move prediction accuracy on the Maia-2 Testset. Skilled, Advanced, and Master are grouped according to Section 4 and Avg denotes macro-averaged results.
@@ 4.1 Move Prediction Accuracy
Maia-2. In Table 1, we show the top-1 move prediction accuracy of all models across all groups of players on the Maia-1 Testset. Maia-2 demonstrates strong and consistent performance across all skill levels, surpassing all baselines. Specifically, despite Maia-1 models being specifically trained to mimic chess moves by players at specific skill levels, Maia-2 emerges as a unified one-for-all model that is consistently effective across the entire spectrum of chess skills. The largest improvement is on Advanced players, where Maia-2 gains 1.5 percentage points over the nearest competitor (Maia 1500). When averaging across skill levels, Maia-2 outperforms all other models by almost 2 full percentage points in overall accuracy. Note that the ceiling accuracy of human move prediction is far below 100% given the randomness and diversity of human decisions—even the same player won’t always make the same decision when faced with the same position. Our 2 percentage point gain is substantial considering that the difference between Maia-1 and Leela, the previous state-of-the-art model for this task and a traditional chess engine not trained for this task at all, is only 6 percentage points. Furthermore, Maia-1 is essentially a mixture of 9 experts targeting the specific players’ skill level, where each expert has 10.3M parameters. Regarding the routing function to select the best-performing expert as a nonparameterized function, Maia-1 has 92M parameters in total. Maia-2, on the other hand, is a one-for-all model with 23.3M parameters under our default settings. Therefore, Maia-2 achieves better human move prediction accuracy with even much fewer trainable parameters.
Baseline models. Both Maia-2 and Maia-1 significantly outperform Stockfish and Leela, typically by 5–15 percentage points. Note that Stockfish and Leela aim to play optimal chess (as most humans do too), and only “predict” human moves when their approximations to optimality happen to overlap with those of human players. However, we compare to these traditional chess engines because besides Maia-1, there are still the default method of creating “human-like” AI agents. The accuracy gap between Maia-1 architectures and traditional chess engines demonstrates the necessity of developing specialized models to mimic human chess moves.
Figure 2: Move prediction accuracy across diverse skill levels. Colors represent performance, with warmer tones indicating higher accuracy.
@@
Move prediction perplexity. While top-1 accuracy gains are important, they may overshadow larger improvements in prediction quality. To account for this, we also measure the perplexity of move predictions , which reflects the model’s confidence in its predictions. A lower perplexity indicates the model is more confident and accurate in human move prediction, as it corresponds to a higher likelihood of the correct human move. As shown in Table 3, Maia-2 consistently yields substantially lower perplexity in all groups of skill levels compared to Maia-1. In particular, Maia-2 significantly outperforms Maia-1 in Advanced and Master moves with relatively large margins, demonstrating the effectiveness of our unified modeling approach across diverse skill levels.
@@
Adaptive move predictions. We now evaluate Maia-2’s ability to predict compare across diverse skill combinations using the Cross-skill Testset. As shown in Figure 2, Maia-2 consistently outperforms both Maia 1100 and Maia 1900 in almost all combinations of active player and opponent player skill levels. In particular, although Maia 1100 and Maia 1900 demonstrate competent performance within their respective domains of expertise, their predictive accuracy decreases substantially outside of these targeted skill levels. This is because Maia-1 models are static and cannot respond to varied skill levels and adjust their predictions accordingly. In contrast, our proposed unified modeling approach with skill-aware attention enables Maia-2 to adapt its predictions to account for the skill levels of both the active player and the opponent player, so that varying skill level configurations correspondingly can result in better aligned human move predictions. More results on comparisons with other Maia-1 versions can be found in Figure 6 in the Appendix.
Move quality. One of our key motivations for creating a unified model of human chess behavior is to guide the development of future algorithmic learning tools. As such, understanding the mistakes that people make is of fundamental interest. Can Maia-2 predict mistakes better than Maia-1? Figure 10 in the Appendix shows the move prediction accuracy on the Grounded Testset as a function of move quality, measured by win-rate loss, which is calculated following the same procedures as prior studies [1, 25]. All models generally decrease in their ability to predict worse moves, since humans are generally trying to avoid mistakes, and high-quality moves are more certain whereas lower-quality moves can be more random and thus hard to predict. Nevertheless, Maia-2 outperforms all versions of Maia-1 across most of the move quality range, demonstrating the effectiveness of our unified modeling approach for human move prediction.
We are also interested in how certain the models are about their predictions of various move qualities. Figure 3.(A)(top) shows the probabilities that Maia-2 (x-axis) and Maia-1 (y-axis) attribute to the moves people actually played in Grounded Testset. The concentration of points in the lower right quadrant (closer to P(Maia-2) = 1 and P(Maia 1900) = 0) suggests that Maia-2 assigns a higher probability to the true move than Maia-1 does, indicating superior predictive performance. Conversely, the less dense upper left quadrant indicates fewer instances where Maia 1900 outperforms Maia-2. Remarkably, while this consistently occurs across all move qualities, the distinction is more pronounced for Blunders and Errors compared to Optimal moves. Additionally, the bottom row of Figure 3.(A) shows the log odds ratio between $P(x,y)$ and $P(y,x)$ in the top row. The abundance of blue points below the diagonal indicates that Maia-2 is almost always more confident in the correct move than Maia-1 is, indicating an across-the-board improvement in move prediction. Maia-2 offers superior and more confident prediction across diverse move qualities. We also conduct an ablation study as shown in Table 3, we point the readers to Appendix A for more information.
Figure 3: (A). (Top) Joint probability assigned to human moves played by Maia-2 ($x$) and Maia 1900 ($y$), split by move quality. Blunders (left) reduce the expected win-rate by $\geq$ 10%, Errors (middle) by 5–10%, and Optimal (right) by $\leq 0\%$. (Bottom) Log odds ratio of $p(x,y)$ and $p(y,x)$ from top. (B). Move prediction agreement as (left) active player and (right) opponent player skill are varied. All cells are evaluated on the same set of positions but with altered skill level configurations.